OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 6, an = 2 • an-1

OpenStudy (anonymous):

@calculusxy

OpenStudy (amistre64):

let n=2 ... whats the result?

OpenStudy (anonymous):

im not really ssure could you walk me through it?:)

OpenStudy (amistre64):

a1 = 6 an = 2 * an-1 ; let n=2 a2 = 2 * a2-1 = 2 * a1 ; and we know a1 = 6 = 2*6 =12 a1 = 6, a2 = 12 let n=3 ...

OpenStudy (anonymous):

ok

OpenStudy (calculusxy):

Okay. So you can use that formulae for this question also. Just replace the numbers to where the variables were.

OpenStudy (anonymous):

I got 6, 12, 24, 48, 96, 192 @amistre64

OpenStudy (amistre64):

that is correct

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

tahnks dude

OpenStudy (amistre64):

youre welcome